OpenStudy (anonymous):

what is the center of this circle? 7x squared plus 7y squared minus 98x minus 112y plus 679=0

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

what is the center of this circle? 7x squared plus 7y squared minus 98x minus 112y plus 679=0

OpenStudy (anonymous):

7x²+7y²-98x-112y+679=0 7(x²+y²-14x-16y)+679=0 7(x-7)²+7(y-8)²=112 So the center is (7,8)

OpenStudy (anonymous):

what is the radius of this equation: 3x squared plus 3y squared plus 12x minus 24y minus 48=0